OpenStudy (countrygirl1431):

Solve the following equation. Then place the correct answer in the box provided. 2x + 1 < 9

OpenStudy (countrygirl1431):

@rishavraj

OpenStudy (jameshorton):

2x - 1 = 9 2x = 9 + 1 2x = 10 (divide both sides by 2 to get x) 2x/2 = 10/2 x = 5